Solve the enthalpy of methane formation using the following reaction equations


  1. "2H\u2082(g)+O\u2082(g) \u2192 2H\u2082O(g)" "\u2206H = -571,6 kJ"
  2. "\ufeffC(s)+O\u2082(g) \u2192 CO\u2082(g)" "\u2206H = -393,5 kJ"
  3. CH4 (g) + 2 O2 (g) → CO2 (g) + 2 H2O (g) "\u2206H = -889,8 kJ"
1
Expert's answer
2021-01-28T05:38:13-0500

eq - equation

(1st eq): 2H2(g) + O2(g) → 2H2O(g), ∆H = -571.6 kJ/mol

(2nd eq): C(s) + O2(g) → CO2(g), ∆H = -393.5 kJ/mol

(3d eq): CH4(g) + 2O2(g) → CO2(g) + 2H2O(g), ∆H = -889.8 kJ/mol


C(s) + 2H2(g) → CH4(g), ∆H = ???


Solution:

According to Hess's law, the heat of reaction depends upon Initial and final conditions of reactants and does not depend of the intermediate path of the reaction.


1) Modify the three given equations to get the target equation:

a) (1st eq): do nothing. We need 2H2(g) on the reactant side and that's what we have.

b) (2nd eq): do nothing. We need one C(s) on the reactant side and that's what we have.

c) (3d eq): flip it so as to put CH4(g) on the product side.


2) Rewrite all three equations with the changes made (including changes in their enthalpy):

(1st eq): 2H2(g) + O2(g) → 2H2O(g), ∆H = -571.6 kJ/mol

(2nd eq): C(s) + O2(g) → CO2(g), ∆H = -393.5 kJ/mol

(3*deq): CO2(g) + 2H2O(g) → CH4(g) + 2O2(g), ∆H = +889.8 kJ/mol


3) Cancel out the common species on both sides:

2O2(g) ⇒ sum of 1st and 2nd equation & 3*d equation

2H2O(g) ⇒ 3*d & 1st equation

CO2(g) ⇒ 3*d & 2nd equation


Thus, adding modified equations and canceling out the common species on both sides, we get:

C(s) + 2H2(g) → CH4(g)


5) Add the ΔH values of (1st), (2nd) and (3*d) equations to get your answer:

∆Hx = (-571.6 kJ/mol) + (-393.5 kJ/mol) + 889.9 kJ/mol = -75.3 kJ/mol

Hence the enthalpy of formation of methane (CH4) is -75.3 kJ/mol.


Answer: The enthalpy of formation of methane (CH4) is -75.3 kJ/mol.